If the moon orbits around the Earth once every 28 days, what would be the moons orbital period if it's mass was reduced to 1/4?
Answer #1

Orbital speed V= sqrt (GM/R)

When M'= M/4

So new velocity, V' = V/2

So new velocity will be half of older velocity. (ans)